A sphere has a great circle with a circumference of 330 centimeters. (a) What is the radius of the sphere? (b) What is the surface area of the sphere? Round the answers to the nearest tenth and label the answers correctly.

Answer by Boreal(15235)

C=2*pi*r
330=2*pi*r
r=330/(2*pi)=52.5 cm.
A=4*pi*r^2=34,663.9 cm^2, using the 52.5 cm above.